OpenStudy (anonymous):

Help Math trouble plz asap win medal Emilia simplified the equation a-3/2=1/6(3a-9) as shown. What do you obsereve? What does this imply about the solution to the original equation? a-3/2=1/6(3a-9) a-3/2=3a/6-9/6 4-3/2=(a/2-3/2) a-3/2=a-3/2

OpenStudy (yolomcswagginsggg):

kk lemme see ;/

OpenStudy (anonymous):

ok

OpenStudy (yolomcswagginsggg):

Hmmm

OpenStudy (yolomcswagginsggg):

Nope i can't help on this one.

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

My guess is the order of operations

OpenStudy (anonymous):

Well let me ask you this, are all the steps done correct, how would you go on to do this?

OpenStudy (anonymous):

well i think they may have reduced it when it was a fraction

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

That's fine but is there anything wrong with the algebra?

OpenStudy (anonymous):

Just a quick question, is that 4 suppose to be there?

OpenStudy (anonymous):

Or did you mean to put a and put 4 by accident :P

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Ok so can you do step 3 to step 4 please.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

they reduced the fraction down until they couldn't apply any more numbers

OpenStudy (anonymous):

\[a-3/2=a/2-3/2 \implies a+a/2 = +3/2-3/2 \implies a+a/2 =0\] this is what I was looking for

OpenStudy (anonymous):

So your final solution should a = 0, not a = a.

OpenStudy (anonymous):

ok because a variable can't = a variable thx

OpenStudy (anonymous):

No, it can be a variable just the last step was wrong in this case, it could have very well been a = a.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx